# Learning Enhanced Ensemble Filters

### v1

@@ 1 Introduction
Filtering—determining the conditional distribution of the state of a partially and noisily observed dynamical system given data collected sequentially in time—constitutes a longstanding computational challenge, particularly when the state is high-dimensional. Sequential data assimilation encompasses a broad range of methods aimed at finding the filtering distribution, or simply estimating the state, in an online fashion. The focus of this work is on the use of sequential data assimilation methods for filtering and state estimation, and in particular the methodologies developed primarily in the geophysical sciences, and weather prediction especially, which are relevant in high dimensions. The overarching goal is to introduce a new framework for the machine learning of data assimilation (DA) algorithms, and specifically ensemble filters, which facilitates the sharing of parameters between implementations with different ensemble sizes; we may, for example, train and deploy the same model with different ensemble sizes. To achieve this we work in the context of mean-field state-space formulations of filtering. The resulting algorithms are based on stochastic dynamical systems for a state variable whose evolution depends on its own law; particle approximation leads to implementable algorithms. We show how the attention mechanism, adapted from the transformer methodology at the heart of large language models, can be used to define probability measure-dependent transport maps that can be approximated by empirical measures using arbitrary ensemble sizes. We refer to the resulting class of neural network architectures as measure neural mappings.
@@ 1.1 Literature Review
The Kalman filter [7] is a foundational sequential data assimilation algorithm for linear systems with Gaussian errors. The extended Kalman filter, using linearization, adapts the Kalman methodology to nonlinear systems, but its direct application to high-dimensional geophysical problems is computationally prohibitive. The ensemble Kalman filter (EnKF) [8, 9] addresses this limitation through a Monte Carlo approach, representing statistics via a limited ensemble of model states. Square-root variants, including the ensemble transform Kalman filter and the ensemble adjustment Kalman filter [10, 11], reduce both storage requirements and sampling errors through deterministic formulations. However, EnKF-based methods commonly encounter challenges in practical implementations and, in particular, suffer from sampling errors caused by finite ensemble sizes. To address these limitations, covariance inflation [12, 13] and localization [14] techniques have been developed. Approaches such as the iterative EnKF [15] and the maximum likelihood ensemble filter [16] further improve performance in nonlinear regimes.
@@
Complementary to these direct learning strategies, a distinct class of approaches formulates ensemble filtering as a transport problem. These methods explicitly construct transformations that map forecast distributions onto analysis distributions. Representative techniques include Knothe–Rosenblatt rearrangements [24], normalizing flows [25], bridge matching [26], and optimal transport frameworks [27]. These methods offer rigorous probabilistic interpretations and enable accurate sampling from the filtering distribution at each assimilation step. Typically they operate at the distributional level and are not designed to directly transform or update the ensemble members collectively.
@@ 1.2 Notation
For a bounded open set $\Omega\in\mathbb{R}^{d_{\Omega}}$, we use notation $\mathcal{U}(\Omega;\mathbb{R}^{d})$ to denote a Banach space of $\mathbb{R}^{d}-$ valued functions defined on $\Omega$. The symbol $\mathcal{U}$ will distinguish different classes of function, for example those of Hölder or Sobolev regularity.
We denote by $\mathcal{P}(\Omega)$ the space of probability measures defined on set $\Omega$. Given a sigma-algebra on $\Omega$, we define a probability measure $\mathbb{P}$ on $\Omega$ and we let $\mathbb{E}$ denote expectation under this probability probability measure. We denote by $\operatorname{Law}(v)$ the law of a random variable $v$, and denote by $T_{\#}\pi$ the pushforward measure of $\pi$ by the map $T$: if $u\sim\pi$ then $T(u)\sim T_{\#}\pi.$ For $\tau\in\mathbb{R}^{d}$, we let $\delta_{\tau}$ denote the Dirac mass on $\mathbb{R}^{d}$ centered at point $\tau$. We use the notation $\mathsf{N}(m,C)$ for the Gaussian distribution with mean $m$ and covariance $C$. We use the font $\mathsf{mathsf}$ for operators acting on the space of probability measures or the space of functions; we also use this convention for operations on the space of vector-valued sequences over $[N]$, denoted $\mathcal{U}([N];\mathbb{R}^{d})$. If an operator $\mathsf{B}$ on probability measures is defined as the pushforward by a map, we denote the associated transport map with the font $\mathfrak{mathfrak}$, i.e. $\mathsf{B}(\cdot)=\mathfrak{B}_{\sharp}(\cdot)$.
In the remainder of this paper we will use assume that all probability measures are absolutely continuous with respect to the Lebesgue measure so that they have probability density functions, or that they comprise a convex combination of Dirac masses. We will refer to probability measures and probability density functions interchangeably, depending on the setting.
@@ 1.3 Contributions and Overview
    We demonstrate that the resulting filtering algorithm outperforms an optimized local ensemble transform Kalman filter (LETKF, a state-of-the-art ensemble filter) on Lorenz ’96, Kuramoto–Sivashinsky, and Lorenz ’63 models, for both small and larger ensemble sizes.
@@ 2.1 Filtering Set-Up
where we assume that the sequence $\{\xi_{j}^{\dagger}\}_{j\in\mathbb{Z}^{+}}$ is independent of initial condition $v_{0}^{\dagger}$; this is often written as $\{\xi_{j}^{\dagger}\}_{j\in\mathbb{Z}^{+}}\perp v_{0}^{\dagger}$. The data model is given by
@@
where we assume that the sequence $\{\eta_{j}^{\dagger}\}_{j\in\mathbb{N}}\perp v_{0}^{\dagger}$ and that the two sequences in the dynamics and data models are independent of one another: $\{\xi_{j}^{\dagger}\}_{j\in\mathbb{Z}^{+}}\perp\{\eta_{j}^{\dagger}\}_{j\in\mathbb{N}}$. The states $v_{j}^{\dagger}$ lie in $\mathbb{R}^{d_{v}}$, while the observations $y_{j}^{\dagger}$ lie in $\mathbb{R}^{d_{y}}$. Note that there is a hidden parameter $\Delta t$ specifying the time interval between steps $j$ and $j+1$. It is implicitly embedded in (2.1) through $\Psi$, which represents the discrete-time map obtained by integrating the continuous dynamics over an interval of length $\Delta t$.
@@ 2.3 Filtering: State-Space Evolution
Predict: $\quad\widehat{v}_{j+1}$ $=\Psi(v_{j})+\xi_{j},$ (2.9a)
Extend to observation space: $\quad\widehat{y}_{j+1}$ $=h(\widehat{v}_{j+1})+\eta_{j+1},$ (2.9b)
Analysis: $v_{j+1}$ $=\mathfrak{B}(\widehat{v}_{j+1},\widehat{y}_{j+1},y^{\dagger}_{j+1},\rho_{j+1}),$ (2.9c)
@@
$\mathsf{B}(\rho_{j+1};y^{\dagger}_{j+1})=\mathfrak{B}(\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,,\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,,;y^{\dagger}_{j+1};\rho_{j+1})_{\sharp}(\rho_{j+1}).$
@@
$v_{j+1}=\mathfrak{B}_{\mathrm{EnKF}}(\widehat{v}_{j+1},\widehat{y}_{j+1},y^{\dagger}_{j+1},\rho_{j+1}),\$ (2.10)
@@
$\widehat{C}_{j+1}^{vh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}\left[\left(\widehat{v}-\bar{v}_{j+1}\right)\otimes\left(h(\widehat{v})-\bar{h}_{j+1}\right)\right],$ (2.12a)
$\widehat{C}_{j+1}^{hh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}\left[\left(h(\widehat{v})-\bar{h}_{j+1}\right)\otimes\left(h(\widehat{v})-\bar{h}_{j+1}\right)\right].$ (2.12b)
Here, $\bar{v}_{j+1}$ and $\bar{h}_{j+1}$ are the mean of states and observations respectively, under $\widehat{\pi}_{j+1}$:
$\bar{v}_{j+1}=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}\widehat{v},\quad\bar{h}_{j+1}=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}h(\widehat{v}).$ (2.13)
@@
$v^{(n)}_{j+1}$ $=\mathfrak{B}_{\mathrm{EnKF}}\bigl{(}\widehat{v}^{(n)}_{j+1},\widehat{y}^{(n)}_{j+1},y^{\dagger}_{j+1},\rho_{j+1}^{(N)}\bigr{)}.$ (2.15c)
@@
$\widehat{C}_{j+1}^{vh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}\left[\left(\widehat{v}-\bar{v}_{j+1}\right)\otimes\left(h(\widehat{v})-\bar{h}_{j+1}\right)\right]$
$=\frac{1}{N}\sum_{n=1}^{N}\left(\widehat{v}^{(n)}_{j+1}-\bar{v}_{j+1}\right)\otimes\left(h(\widehat{v}^{(n)}_{j+1})-\bar{h}_{j+1}\right),$ (2.16a)
$\widehat{C}_{j+1}^{hh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}\left[\left(h(\widehat{v})-\bar{h}_{j+1}\right)\otimes\left(h(\widehat{v})-\bar{h}_{j+1}\right)\right]$
$=\frac{1}{N}\sum_{n=1}^{N}\left(h(\widehat{v}^{(n)}_{j+1})-\bar{h}_{j+1}\right)\otimes\left(h(\widehat{v}^{(n)}_{j+1})-\bar{h}_{j+1}\right).$ (2.16b)
Here $\bar{v}_{j+1}$ and $\bar{h}_{j+1}$, originally defined in (2.13), are also computed by use of an empirical approximation of $\widehat{\pi}_{j+1}$:
$\bar{v}_{j+1}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}\widehat{v}=\frac{1}{N}\sum_{n=1}^{N}\widehat{v}^{(n)}_{j+1},$ (2.17a)
$\bar{h}_{j+1}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}h(\widehat{v})=\frac{1}{N}\sum_{n=1}^{N}h(\widehat{v}^{(n)}_{j+1}).$ (2.17b)
@@
The ensemble Kalman filter exhibits permutation invariance with respect to the ensemble members. This important property stems from the fact that both the sample means $\bar{v}_{j+1}$, $\bar{h}_{j+1}$ in (2.17) and the empirical covariance matrices $\widehat{C}_{j+1}^{vh}$, $\widehat{C}_{j+1}^{hh}$ in (2.16) are calculated as averages over all ensemble members, making them invariant to the ordering of particles in the ensemble. This permutation invariance is a crucial property that motivates the design of our methodology.
@@ 2.4 Learning Probability Measure-Dependent Analysis Maps
$\mathfrak{B}_{\textrm{MNM}}(\widehat{v}_{j+1},\widehat{y}_{j+1},y^{\dagger}_{j+1},\rho_{j+1})\approx\mathfrak{B}(\widehat{v}_{j+1},\widehat{y}_{j+1},y^{\dagger}_{j+1},\rho_{j+1}),$ (2.19a)
$\mathsf{B}_{\textrm{MNM}}(\rho_{j+1};y_{j+1}^{\dagger})=\mathfrak{B}_{\textrm{MNM}}(\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,,\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,,y^{\dagger}_{j+1},\rho_{j+1})_{\sharp}\rho_{j+1}\approx\mathsf{B}(\rho_{j+1};y_{j+1}^{\dagger}).$ (2.19b)
@@
$v_{j+1}=\mathfrak{B}_{\textrm{MNM}}(\widehat{v}_{j+1},\widehat{y}_{j+1},y^{\dagger}_{j+1},\rho_{j+1}).$ (2.20)

### v2

@@ Abstract
{highlights}
We demonstrate that the resulting filtering algorithm outperforms an optimized local ensemble transform Kalman filter (LETKF, a leading ensemble filter) on Lorenz ’96, Kuramoto–Sivashinsky, and Lorenz ’63 models, for both small and larger ensemble sizes.
@@ 1 Introduction
Filtering—determining the conditional distribution of the state of a partially and noisily observed dynamical system given data collected sequentially in time—constitutes a longstanding computational challenge, particularly when the state is high-dimensional. Sequential data assimilation encompasses a broad range of methods aimed at finding the filtering distribution, or simply estimating the state, in an online fashion. The focus of this work is on the use of sequential data assimilation methods for filtering and state estimation, and in particular the methodologies developed primarily in the geophysical sciences, and weather prediction especially, which are relevant in high dimensions. The overarching goal is to introduce a new framework for the machine learning of data assimilation (DA) algorithms, and specifically ensemble filters, which facilitates the sharing of parameters between implementations with different ensemble sizes; we may, for example, train and deploy the same model with different ensemble sizes. To achieve this we work in the context of mean-field state-space formulations of filtering. The resulting algorithms are based on stochastic dynamical systems for a state variable whose evolution depends on its own law; particle approximation leads to implementable algorithms. We show how the attention mechanism, adapted from the transformer methodology at the heart of large language models, can be used to define probability measure-dependent transport maps, i.e. mappings that move probability mass from one distribution to another, that can be approximated by empirical measures using arbitrary ensemble sizes. We refer to the resulting class of neural network architectures as measure neural mappings.
@@ 1.1 Literature Review
The Kalman filter [7] is a foundational sequential data assimilation algorithm for linear systems with Gaussian errors. The extended Kalman filter, using linearization, adapts the Kalman methodology to nonlinear systems, but its direct application to high-dimensional geophysical problems is computationally prohibitive. The ensemble Kalman filter (EnKF) [8, 9, 10] addresses this limitation through a Monte Carlo approach, representing statistics via a limited ensemble of model states. Square-root variants, including the ensemble transform Kalman filter and the ensemble adjustment Kalman filter [10, 11], reduce both storage requirements and sampling errors through deterministic formulations. However, EnKF-based methods commonly encounter challenges in practical implementations and, in particular, suffer from sampling errors caused by finite ensemble sizes. To address these limitations, covariance inflation [12, 13] and localization [14] techniques have been developed. Approaches such as the iterative EnKF [15] and the maximum likelihood ensemble filter [16] further improve performance in nonlinear regimes.
@@
Complementary to these direct learning strategies, a distinct class of approaches formulates ensemble filtering as a transport problem. These methods explicitly construct transformations that map forecast distributions onto analysis distributions. Representative techniques include Knothe–Rosenblatt rearrangements [24], normalizing flows [25], bridge matching [26], and optimal transport frameworks [27]. These methods offer rigorous probabilistic interpretations and enable accurate sampling from the filtering distribution at each assimilation step. Typically these approaches find transformations acting on each individual particle, rather than operating directly at the distributional level, and are not designed to directly transform or update the ensemble members collectively.
@@ 1.2 Notation
We denote by $\mathcal{P}(\Omega)$ the space of probability measures defined on set $\Omega$. Given a sigma-algebra on $\Omega$, we define a probability measure $\mathbb{P}$ on $\Omega$ and we let $\mathbb{E}$ denote expectation under this probability measure. We denote by $\operatorname{Law}(v)$ the law of a random variable $v$, and denote by $T_{\#}\pi$ the pushforward measure of $\pi$ by the map $T$: if $u\sim\pi$ then $T(u)\sim T_{\#}\pi.$ For $\tau\in\mathbb{R}^{d}$, we let $\delta_{\tau}$ denote the Dirac mass on $\mathbb{R}^{d}$ centered at point $\tau$. We use the notation $\mathsf{N}(m,C)$ for the Gaussian distribution with mean $m$ and covariance $C$. We use the font $\mathsf{mathsf}$ for operators acting on the space of probability measures or the space of functions; we also use this convention for operations on the space of vector-valued sequences over $[N]$, denoted by $\mathcal{U}([N];\mathbb{R}^{d})$. If an operator $\mathsf{B}$ on probability measures is defined as the pushforward by a map, we denote the associated transport map with the font $\mathfrak{mathfrak}$, i.e. $\mathsf{B}(\cdot)=\mathfrak{B}_{\sharp}(\cdot)$.
In the remainder of this paper we assume that all probability measures are absolutely continuous with respect to the Lebesgue measure so that they have probability density functions, or that they comprise a convex combination of Dirac masses. We will refer to probability measures and probability density functions interchangeably, depending on the setting.
@@ 1.3 Contributions and Overview
    We demonstrate that the resulting filtering algorithm outperforms an optimized local ensemble transform Kalman filter (LETKF, a leading ensemble filter) on Lorenz ’96, Kuramoto–Sivashinsky, and Lorenz ’63 models, for both small and larger ensemble sizes.
@@ 2.1 Filtering Set-Up
where $\Psi:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{v}}$ is the forward dynamic model, and $v^{\dagger}_{j},\xi_{j}^{\dagger}\in\mathbb{R}^{d_{v}}$ for $j\in\mathbb{Z}^{+}$. We assume that the sequence $\{\xi_{j}^{\dagger}\}_{j\in\mathbb{Z}^{+}}$ is independent of initial condition $v_{0}^{\dagger}$; this is often written as $\{\xi_{j}^{\dagger}\}_{j\in\mathbb{Z}^{+}}\perp v_{0}^{\dagger}$. The data model is given by
@@
where $h:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{y}}$ is the observation operator and $y^{\dagger}_{j},\eta_{j}^{\dagger}\in\mathbb{R}^{d_{y}}$ for $j\in\mathbb{Z}^{+}$. We assume that the sequence $\{\eta_{j}^{\dagger}\}_{j\in\mathbb{N}}\perp v_{0}^{\dagger}$ and that the two sequences in the dynamics and data models are independent of one another: $\{\xi_{j}^{\dagger}\}_{j\in\mathbb{Z}^{+}}\perp\{\eta_{j}^{\dagger}\}_{j\in\mathbb{N}}$. The states $v_{j}^{\dagger}$ lie in $\mathbb{R}^{d_{v}}$, while the observations $y_{j}^{\dagger}$ lie in $\mathbb{R}^{d_{y}}$. Note that there is a hidden parameter $\Delta t$ specifying the time interval between steps $j$ and $j+1$. It is implicitly embedded in (2.1) through $\Psi$, which represents the discrete-time map obtained by integrating the continuous dynamics over an interval of length $\Delta t$.
@@ 2.3 Filtering: State-Space Evolution
Predict: $\quad\widehat{v}_{j+1}$ $=\Psi(v_{j})+\xi_{j},\quad\xi_{j}\sim\mathsf{N}(0,\Sigma),$ (2.9a)
Extend to observation: $\quad\widehat{y}_{j+1}$ $=h(\widehat{v}_{j+1})+\eta_{j+1},\quad\eta_{j+1}\sim\mathsf{N}(0,\Gamma),$ (2.9b)
Analysis: $v_{j+1}$ $=\mathfrak{B}(\widehat{v}_{j+1},\widehat{y}_{j+1};y^{\dagger}_{j+1},\rho_{j+1}),$ (2.9c)
@@
$\mathsf{B}(\rho_{j+1};y^{\dagger}_{j+1})=\mathfrak{B}(\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,,\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,,;y^{\dagger}_{j+1},\rho_{j+1})_{\sharp}(\rho_{j+1}).$
@@
$v_{j+1}=\mathfrak{B}_{\mathrm{EnKF}}(\widehat{v}_{j+1},\widehat{y}_{j+1};y^{\dagger}_{j+1},\rho_{j+1}),\$ (2.10)
@@
$\widehat{C}_{j+1}^{vh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}\left[\left(\widehat{v}-\overline{v}_{j+1}\right)\otimes\left(h(\widehat{v})-\overline{h}_{j+1}\right)\right],$ (2.12a)
$\widehat{C}_{j+1}^{hh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}\left[\left(h(\widehat{v})-\overline{h}_{j+1}\right)\otimes\left(h(\widehat{v})-\overline{h}_{j+1}\right)\right].$ (2.12b)
Here, $\overline{v}_{j+1}$ and $\overline{h}_{j+1}$ are the mean of states and observations respectively, under $\widehat{\pi}_{j+1}$:
$\overline{v}_{j+1}=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}\widehat{v},\quad\overline{h}_{j+1}=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}_{j+1}}h(\widehat{v}).$ (2.13)
@@
$v^{(n)}_{j+1}$ $=\mathfrak{B}_{\mathrm{EnKF}}\bigl{(}\widehat{v}^{(n)}_{j+1},\widehat{y}^{(n)}_{j+1};y^{\dagger}_{j+1},\rho_{j+1}^{(N)}\bigr{)}.$ (2.15c)
@@
$\widehat{C}_{j+1}^{vh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}\left[\left(\widehat{v}-\overline{v}_{j+1}\right)\otimes\left(h(\widehat{v})-\overline{h}_{j+1}\right)\right]=\frac{1}{N}\sum_{n=1}^{N}\left(\widehat{v}^{(n)}_{j+1}-\overline{v}_{j+1}\right)\otimes\left(h(\widehat{v}^{(n)}_{j+1})-\overline{h}_{j+1}\right),$ (2.16a)
$\widehat{C}_{j+1}^{hh}$ $=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}\left[\left(h(\widehat{v})-\overline{h}_{j+1}\right)\otimes\left(h(\widehat{v})-\overline{h}_{j+1}\right)\right]=\frac{1}{N}\sum_{n=1}^{N}\left(h(\widehat{v}^{(n)}_{j+1})-\overline{h}_{j+1}\right)\otimes\left(h(\widehat{v}^{(n)}_{j+1})-\overline{h}_{j+1}\right).$ (2.16b)
Here $\overline{v}_{j+1}$ and $\overline{h}_{j+1}$, originally defined in (2.13), are also computed by use of an empirical approximation of $\widehat{\pi}_{j+1}$:
$\overline{v}_{j+1}=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}\widehat{v}=\frac{1}{N}\sum_{n=1}^{N}\widehat{v}^{(n)}_{j+1},\qquad\overline{h}_{j+1}=\mathbb{E}^{\widehat{v}\sim\widehat{\pi}^{(N)}_{j+1}}h(\widehat{v})=\frac{1}{N}\sum_{n=1}^{N}h(\widehat{v}^{(n)}_{j+1}).$ (2.17a)
@@
The ensemble Kalman filter exhibits permutation invariance with respect to the ensemble members. This important property stems from the fact that both the sample means $\overline{v}_{j+1}$, $\overline{h}_{j+1}$ in (2.17) and the empirical covariance matrices $\widehat{C}_{j+1}^{vh}$, $\widehat{C}_{j+1}^{hh}$ in (2.16) are calculated as averages over all ensemble members, making them invariant to the ordering of particles in the ensemble. This permutation invariance is a crucial property that motivates the design of our methodology.
@@ 2.4 Learning Probability Measure-Dependent Analysis Maps
$\mathfrak{B}_{\textrm{MNM}}(\widehat{v}_{j+1},\widehat{y}_{j+1};y^{\dagger}_{j+1},\rho_{j+1})\approx\mathfrak{B}(\widehat{v}_{j+1},\widehat{y}_{j+1};y^{\dagger}_{j+1},\rho_{j+1}),$ (2.19a)
$\mathsf{B}_{\textrm{MNM}}(\rho_{j+1};y_{j+1}^{\dagger})=\mathfrak{B}_{\textrm{MNM}}(\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,,\,\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet}\,;y^{\dagger}_{j+1},\rho_{j+1})_{\sharp}\rho_{j+1}\approx\mathsf{B}(\rho_{j+1};y_{j+1}^{\dagger}).$ (2.19b)
@@
$v_{j+1}=\mathfrak{B}_{\textrm{MNM}}(\widehat{v}_{j+1},\widehat{y}_{j+1};y^{\dagger}_{j+1},\rho_{j+1}).$ (2.20)